OpenStudy (anonymous):

Can you guys please help and check my answers for me? To see if I got them right.

OpenStudy (anonymous):

1. What is the y-intercept of the graph of y = 3x2 6x 8? (Points : 4) -8 6 1 3 I think it is A for this one.

OpenStudy (bibby):

the y intercept is the point on the graph where x=0 replace x with 0, you get -8, I presume

OpenStudy (anonymous):

I think this one is also A

OpenStudy (bibby):

the axis of symmetry formula is \(\large x=\frac{-b}{2a}\)

OpenStudy (anonymous):

So my answer is wrong?

OpenStudy (bibby):

I forgot to check

OpenStudy (bibby):

looks good, sorry

OpenStudy (anonymous):

Yes the answer is incorrect.

OpenStudy (anonymous):

Umm.? Ha. now I'm confused.

OpenStudy (bibby):

\(b=-5\\ a=-2\\ 2a=-4\\ \large \frac{-b}{2a}=\frac{--5}{-4}=\frac{-5}{4}\)

OpenStudy (anonymous):

So I'm right.?

OpenStudy (anonymous):

A again for this one

OpenStudy (bibby):

that's what I said, yeah

OpenStudy (bibby):

start by finding x with the same formula as the previous problem. I got x=2. then plug it into the equation, I got 4-8-5=4-13=-9

OpenStudy (bibby):

correct

OpenStudy (anonymous):

I chose that graph as the answer

OpenStudy (anonymous):

That's what she said.

OpenStudy (bibby):

to do this kind of problem, find the x intercepts by setting y to 0. we then know that \(a \times b=0\) when either a or b is 0. when a is 0: \(x+1=0\longrightarrow x=-1\) when b is 0: \(x-4=0\longrightarrow x=4\)

OpenStudy (anonymous):

So that's the wrong graph

OpenStudy (bibby):

mhm

OpenStudy (bibby):

my friend is sending me snapchats of himself pissing, send help

OpenStudy (anonymous):

So its this one?

OpenStudy (bibby):

-1 and 4 not -1 and -4

OpenStudy (anonymous):

You should send me some, too.

OpenStudy (anonymous):

Okay one ew xD and two I'm confused. So would it be this one then because the others look really wrong

OpenStudy (bibby):

add me, bibums

OpenStudy (bibby):

as I understand it, it'd look something like